OpenStudy (anonymous):

Solve the system, using substitution. What is the x coordinate of the solution? x = 2y - 2 and 3x - y = 14 A. 6 B. 4 C. 1 D. 0 E. no solution

OpenStudy (anonymous):

whats the answer

jimthompson5910 (jim_thompson5910):

3x - y = 14 3(2y-2) - y = 14 ... replace x with 2y - 2 6y - 6 - y = 14 5y - 6 = 14 I'll let you finish up

OpenStudy (anonymous):

i dont know how to do this

OpenStudy (anonymous):

so one solution?

jimthompson5910 (jim_thompson5910):

are you able to solve 5y - 6 = 14 ?

OpenStudy (anonymous):

yeah so its a?

jimthompson5910 (jim_thompson5910):

what solution do you get for y?

OpenStudy (anonymous):

how did he do that?

OpenStudy (anonymous):

you*

jimthompson5910 (jim_thompson5910):

which part?

OpenStudy (anonymous):

The solving part and was 4 the answer?

jimthompson5910 (jim_thompson5910):

have a look at this article http://www.purplemath.com/modules/solvelin.htm be sure to go through the 4 pages